OpenStudy (anonymous):

I need help integrating x/x^2y+4y

geerky42 (geerky42):

With respect to what?

OpenStudy (anonymous):

dx

OpenStudy (anonymous):

Set \(u=y(x^2+4)\), then \(du=2xy~dx\), i.e. \(\dfrac{1}{2y}~du=x~dx\). \[\int\frac{x}{y(x^2+4)}~dx=\frac{1}{2y}\int\frac{du}{u}\]